# ZeroSearch: Incentivize the Search Capability of LLMs without Searching

### v1

@@ 5.3 Interaction Turns Study
PPO 39.92 62.15 43.61 29.12 35.44 12.98 15.32 34.08
GRPO 40.28 61.87 49.80 30.57 35.05 14.75 14.52 35.26
Table 5: Performance of ZeroSearch under different RL algorithms. We compare PPO and GRPO using the Qwen2.5-3B-Base and LLaMA-3.2-3B-Base models.
In this section, we evaluate the performance of two widely adopted RL training algorithms, PPO and GRPO, within the ZeroSearch framework, using the Qwen2.5-3B-Base and LLaMA-3.2-3B-Base models. The results of this comparison are presented in Section 5.4.
As observed, both GRPO and PPO successfully incentivize the search capabilities within our framework, demonstrating the versatility of our approach. Among them, GRPO exhibits more stable performance across both models, emphasizing its advantage in terms of training stability. It is also worth noting that the repeated rollout mechanism in GRPO incurs higher API costs when interacting with a real search engine, further highlighting the practicality of our simulated search setting.
5.5 Reverse Curriculum Study
Method NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle Avg.
Qwen-2.5-3B-Base
Easy $\rightarrow$ Hard 41.21 61.49 43.99 31.02 33.20 12.58 14.29 33.97
Hard $\rightarrow$ Easy 39.59 58.78 44.31 32.32 30.67 12.55 12.00 32.89
[…]
Easy $\rightarrow$ Hard 66.19 57.29 23.39 24.44 30.75 9.39 19.67 33.02
Hard $\rightarrow$ Easy 40.16 57.83 24.69 28.72 27.55 8.96 8.94 28.12
Table 6: Reverse Curriculum Study. We compare the performance of standard and reverse curriculum rollout settings using the Qwen-2.5-3B-Base and Qwen-2.5-3B-Instruct models.
In this section, we analyze the effectiveness of the curriculum rollout strategy by comparing it with a reverse curriculum setup, in which the difficulty of training decreases over time by gradually improving the quality of retrieved documents. The results are presented in Section 5.5.
The results clearly indicate that the standard easy-to-hard curriculum consistently outperforms the reverse hard-to-easy variant across both models, supporting the effectiveness of curriculum learning in our framework. Starting with better search results allows the policy model to first learn how to invoke the search engine and understand the basic output format. As training progresses, the model is exposed to increasingly challenging scenarios, fostering stronger reasoning capabilities.
6 Conclusion
In this paper, we propose ZeroSearch, a novel RL framework that enhances the search capabilities of LLMs without interacting with real search engines. Through supervised fine-tuning, the LLM is transformed into a retrieval module capable of generating both relevant and noisy documents. A curriculum rollout mechanism is employed to progressively improve reasoning by exposing the model to increasingly challenging retrieval scenarios. Experimental results show that ZeroSearch outperforms real search-based models, generalizes well across both base and instruction-tuned LLMs of varying sizes, and supports a wide range of RL algorithms.
However, our approach has certain limitations. Deploying the simulated search LLM requires access to GPU servers. While more cost-effective than commercial API usage, this introduces additional infrastructure costs. We provide a detailed discussion of these costs in the appendix.
@@ Appendix A Compare ZeroSearch with Real Search Engine
In this section, we present additional results comparing ZeroSearch with a real search engine using the Qwen-2.5-3B and Qwen-2.5-7B model series in Figure 3.
@@ Appendix E Case Study
Question: Who is the spouse of the person who does the voice for Smokey the Bear?
@@
<think> I need to find the voice actor for Smokey the Bear and then find his spouse. </think> <search> Smokey the Bear voice actor </search> <information> Doc 1: Stephen Colbert, Al Roker and Jeff Foxworthy have joined the historic campaign, lending their voices to help expand on Smokey’s iconic catchphrase. Doc 2: TIL actor Sam Elliott – known for his full mustache, and deep, resonant voice – voices Smokey the Bear in the iconic character’s public Doc 3: Actor Sam Elliott, on sharing a birthday and voice with Smokey Bear With his bright, white mustache and earthy, western drawl, Sam Elliott Doc 4: Sam Elliot was a great Smokey Bear! Other actors who probably could have voiced him include Kevin Tighe and Randolph Mantooth whose character on Doc 5: Shares birthday with "Smokey Bear," and Sam Elliott has been the voice of "Smokey Bear" since 2008. Inducted into the International Mustache Hall of Fame in </information> <think> The voice actor for Smokey the Bear is Sam Elliott. Now I need to find his spouse. </think> <search> Sam Elliott spouse </search> <information> Doc 1: Sam Elliot and Katharine Ross have been married for 35 years. Here’s how they keep the spark alive even when working together on a film. Doc 2: Meet Sam Elliott’s And Katharine Ross’ 40-Year Love Story#hollywood · Comments1.8K. Doc 3: Sam Elliott and Katharine Ross met in 1978 and were married in 1984. They met while filming the horror movie, The Legacy. It was Ross’s fifth Doc 4: Sam Elliott and his wife Katharine Ross have been married for nearly four decades. Their love story dates back to the late 1960s. Doc 5: 1978 Pete (Sam Elliott) and Maggie (Katherine Ross) are lovers to whom strange things begin to happen, in Universal’s horror thriller ’The Legacy’ </information> <think> Sam Elliott’s spouse is Katharine Ross. </think> <answer> Katharine Ross </answer>
Question: What is the official color of the third oldest surviving university?
Answer: Cambridge Blue
ZeroSearch:
<think> I need to find the third oldest surviving university and its official color. </think> <search> third oldest surviving university </search> <information> Doc 1: Founded in 1209, Cambridge is the world’s third oldest surviving university and one of its most prestigious, currently ranked second best in the Doc 2: The oldest university in the English-speaking world, Oxford is regularly recognised as one of the premier academic institutions on the planet. Doc 3: The oldest education institution of any kind is likely The King’s School, Canterbury, which has been in continuous operation since 597. Doc 4: Founded in 1088, the University of Bologna is the oldest surviving university in Europe and one of the most prestigious universities in Italy. Doc 5: The University of Oxford is the third oldest university in the world, established in the twelfth century in Oxford, England. With an esteemed </information> <think> The third oldest surviving university is Cambridge University. Now I need to find its official color. </think> <search> official color of cambridge university </search> <information> Doc 1: The logo only appears in the four colour variants shown on this page. Black, Red Pantone 032, Yellow Pantone 109 and white coat of arms with Doc 2: Official University of Cambridge Large Crest Print T-Shirt – Green. £19.99. T-shirt with a one-colour print crest and Cambridge University design. Part of the Doc 3: “The colour is called Cambridge Blue.” “It doesn’t matter what it’s called. It LOOKS green so – it IS green,” she added. I was perplexed and Doc 4: Premium quality unisex relaxed-fit t-shirt with detailed neck-tape and label, embroidered with the University of Cambridge crest. Discontinued colour Doc 5: Cambridge Blue is a soft, muted shade of blue-green that exudes a sense of calm and sophistication. This distinctive color is named after the University of </information> <think> The official color of Cambridge University is Cambridge Blue. </think> <answer> Cambridge Blue </answer>
@@
Question: When was the director of film Reykjavík-Rotterdam born?
Answer: 30 June 1963
Query: Óskar Jónasson birth date
@@
Doc 1: Óskar Jónasson is a television and film director from Iceland. His feature debut Reykjavík-Rotterdam won the Golden Bear at the Berlin International Film Festival
Doc 2: Director of Reykjavík-Rotterdam and Silence of My Father Óskar Jónasson was born on 30 June 1963 in Reykjavík, Iceland. He is a graduate of the Iceland Academy of
Doc 3: A documentary about Iceland’s aborted attempt to build a nuclear power plant in the 1970s, as a result of which the nation became largely
Doc 4: Director of Reykjavík-Rotterdam and Silence of My Father Óskar Jónasson was born on 30 June 1963 in Reykjavík, Iceland. He is a graduate of the Iceland Academy
Doc 5: Icelandic filmmaker Oscar Jonasson talks about his new film, the Golden Bear-winning "Reykjavik-Rotterdam," at the Berlinale this year.
@@
Doc 1: The first single, “Láta ég allt” was released on March 18th 2022. It was written by him along with Óskar Jónasson and is accompanied by music video by the multi-
Doc 2: The children of the first generation include the likes of Ásgeir Trausti (born 1984) and Amiina (formed 2005), who also introduced the world
Doc 3: Óskar Jónasson · 2022 · Iceland · Tracks: 13 · Genres: Pop · Languages: Icelandic.
Doc 4: Óskar Jónasson. · Location. Reykjavík, Iceland. · Industry. Music. · Looking for. Jobs. · Experience. 7 years. · Education. 4 years.
Doc 5: Jónasson was born in Iceland, and his family moved to Sweden when he was a child. He graduated from Stockholm University with a degree in
Question: What is the official color of the third oldest surviving university?
Answer: Cambridge Blue
Query: third oldest surviving university
@@
Doc 1: The University of Oxford is the oldest university in the English-speaking world, with origins that can be traced back to 1096.
Doc 2: The University of Cambridge is the third oldest surviving university in the world, after the University of Bologna in Italy and the University of Paris in
Doc 3: University of Cambridge. The University of Cambridge is the third oldest surviving university in the world, after the University of Bologna in Italy and the
Doc 4: The University of Oxford is the oldest university in the English-speaking world, with origins that can be traced back to 1096.
Doc 5: The University of Cambridge is the third oldest surviving university in the world, after the University of Bologna in Italy and the University of Paris in
@@
Doc 1: The oldest continuously operating university in the world is the University of Bologna, Italy. Founded in 1088, it is the oldest of the
Doc 2: Founded in 1088, the University of Bologna is the oldest in continuous operation. It is also among the most respected universities in the world.
Doc 3: The University of Bologna is the oldest in the world, founded in 1088, and the University of Oxford and the University of Cambridge are both
Doc 4: The University of Bologna is the oldest surviving university in the world, founded in 1088. It is also one of the most prestigious, having educated some of the
Doc 5: Founded in 1088, the University of Bologna is the oldest still in operation. The University of Paris, founded around 1150, is the second oldest.

### v2

@@ 5.5 Curriculum Rollout Study
Curriculum 43.40 63.80 48.40 32.20 35.60 13.80 15.28 36.07
Random 40.40 62.80 49.60 29.80 36.00 14.20 11.11 34.84
Table 6: Curriculum Rollout Study. We compare the performance of standard and random rollout settings using the Qwen-2.5-3B-Base and LLaMA-3.2-3B-Base as the policy models.
In this section, we investigate the effectiveness of our curriculum rollout strategy by comparing it against a baseline random setup, where the probability of generating noisy documents is fixed at 0.5 throughout training. The comparison results are summarized in Section 5.5.
The results clearly indicate that the standard easy-to-hard curriculum consistently outperforms the random rollout variant across both models, verifying the effectiveness of curriculum rollout in our framework. Starting with better search results allows the policy model to first learn how to invoke the search engine and understand the basic output format. As training progresses, the model is exposed to increasingly challenging scenarios, fostering stronger reasoning capabilities.
6 Conclusion and Limitation Discussion
In this paper, we propose ZeroSearch, a novel RL framework that enhances the search capabilities of LLMs without interacting with real search engines. Through supervised fine-tuning, the LLM is transformed into a retrieval module capable of generating both useful and noisy documents. A curriculum rollout mechanism is employed to progressively improve reasoning by exposing the model to increasingly challenging retrieval scenarios. Experimental results show that ZeroSearch outperforms real search-based models, generalizes well across both base and instruction-tuned LLMs of varying sizes, and supports a wide range of RL algorithms.
However, our approach has certain limitations. Deploying the simulated search LLM requires access to GPU servers. While more cost-effective than commercial API usage, this introduces additional infrastructure costs. We provide a detailed discussion of these costs in Appendix D.
@@ Appendix A Compare ZeroSearch with Real Search Engine
In this section, we present additional results comparing ZeroSearch with a real search engine using the LLaMA-3.2-3B and Qwen-2.5-7B model series in Figure 3.
@@ Appendix E Case Study
Question: What is the place of birth of Clara Novello’s father?
Answer: London, England
ZeroSearch:
<think> The question is about the place of birth of Clara Novello’s father. </think> <search> place of birth of Clara Novello’s father </search> <information> Doc 1: 68 Clara Novello’s Reminiscences In this interesting small place I knew the daughter-in-law of Goethe, who showed us his house, his bedroom, the gold laurel Doc 2: So, we may happily add James Clerk Maxwell to our list as this year is the 175th since his birth in Edinburgh on November 13th, 1831. Doc 3: When I was intro- duced to him as Vincent Novello’s eldest daughter, Coleridge was struck by my father’s name, knowing it to be that of a musician, and Doc 4: ABSTRACT. This publication presents biographical profiles of people of interest to young readers. The concept is unique in that. Doc 5: Reuse: Copies of full items can be used for personal research or study, educational, or not-for-profit purposes without prior permission or charge. </information> <think> Based on the search results, the place of birth of Clara Novello’s father is not directly mentioned. However, I can infer that Clara Novello’s father is Vincent Novello, and he is a musician. Since the search results do not provide the place of birth of Vincent Novello, I need to search for his place of birth. </think> <search> place of birth of Vincent Novello </search> <information> Doc 1: Born: September 6, 1781, London, England. Died: August 9, 1861, Nice, France. Buried: Cimetiére du Château, Nice, France. portrait. Biography. Doc 2: NOVELLO, VINCENT (1781–1861), organist, musical composer, editor, and arranger, was born at 240 Oxford Road (now Oxford Street), London, Doc 3: Vincent Novello Leigh Hunt was born about 1823 in Genoa, Italy, where his parents, writer Leigh Hunt and his wife Marianne Kent, had been living Doc 4: NOVELLO, Vincent, son of an Italian father and English mother, was born at 240, Oxford Street, Sept. 6, 1781. Doc 5: Vincent Novello (6 September 1781 – 9 August 1861), was an English musician and music publisher born in London. </information> <think> Based on the search results, the place of birth of Vincent Novello, Clara Novello’s father, is London, England. </think> <answer> London, England </answer>
Question: Who is the spouse of the person who played the sergeant major in We Were Soldiers?
@@
<think> The question is about the spouse of the person who played the sergeant major in the movie "We Were Soldiers". </think> <search> who played the sergeant major in We Were Soldiers </search> <information> Doc 1: Maj. Basil L. Plumley. Sam Elliott was the only actor who Randall Wallace could envision as the battalion’s hard-as-nails sergeant major. Doc 2: But apparently, Plumley had no issues with Elliott’s portrayal. The actor "under played" the sergeant major, said Galloway, who introduced the Doc 3: The U.S. Army is investigating allegations that the late Command Sgt. Major Basil Plumley wore unauthorized combat and valor awards. Doc 4: Plumley was played by actor Sam Elliott in the movie We Were Soldiers (2002) about the Battle of Ia Drang. Command Sergeant Major Basil L. Doc 5: I’ll start by saying Plumley was a badass (edit: apparently not, TIL), and that Sam Elliott’s portrayal was equally brilliant. </information> <think> Based on the search results, the person who played the sergeant major in "We Were Soldiers" is Sam Elliott. Now I need to find out who his spouse is. </think> <search> spouse of Sam Elliott </search> <information> Doc 1: Sam Elliott has been married to Katharine Ross since 1984, indicating a long-term heterosexual relationship. There is no credible evidence to Doc 2: Actor Sam Elliott, star of the film "The Hero," talks Hollywood, his 30-year marriage to actress Katherine Ross and his famous mustache. Doc 3: Sam Elliott and Katharine Ross met in 1978 and were married in 1984. They met while filming the horror movie, The Legacy. It was Ross’s Doc 4: After more than 35 years of marriage, it’s clear that Sam Elliott and his wife, Katharine Ross, are one of Hollywood’s most successful love Doc 5: Sam Elliott and his wife Katharine Ross have been married for nearly four decades. Their love story dates back to the late 1960s. </information> <think> Based on the search results, the spouse of Sam Elliott, who played the sergeant major in "We Were Soldiers," is Katharine Ross. </think> <answer> Katharine Ross </answer>
@@
Question: who came second in tour de france 2018?
Answer: Tom Dumoulin
Query: Tour de France 2018 second place
@@
Doc 1: The 2018 Tour de France was the 105th edition of the race, and the first under the leadership of new event president Étienne Thobois.
Doc 2: 2nd – Tom Dumoulin (NED / Team Sunweb) 3rd – Geraint Thomas (GBR / Team Sky) 4th – Jakob Fuglsang (DEN / Astana) 5th – Romain Bardet (FRA / AG2R
Doc 3: 2018 Tour de France – Stage 20. 15:20 | 51:58. Stage 20 is the final time trial of this year’s Tour de France. The 28.6km parcours will start and finish
Doc 4: Tour de France 2018: Live coverage of final stage | Cycling News · Stage 21 live commentary. 1.1M views · 1 year ago #cyclingnews · Tour de France 2018: Stage 21 -
Doc 5: 2018 TDF. Egan Bernal. 29. Chris Froome. 30. Tom Dumoulin. 31. Geraint Thomas. 32. Romain Bardet. 33. Simon Yates. 34. Thibaut Pinot. 35.
@@
Doc 1: second place is 5000 euros, with 3000 euros for the third-placed rider. Riders in the top three of each stage receive a cash bonus as well.
Doc 2: second place was 1 minute and 47 seconds. This time gap is the biggest ever in the history of Tour de France. The previous record is 1 minute
Doc 3: 1978. Second place, Joop Zoetemelk. 1980. Not Having Enough Time. An hour before Wouter Weylandt won stage nine, a picture of the Tour 2020 got painted.
Doc 4: second place in January 2017. Over the following two years, the Swiss rider won the Giro d’Italia and Tour de France, adding the Vuelta to Spain in
Doc 5: The awards for best riders in their respective categories were presented at the end of each stage and the overall rankings are published every evening.
Question: What is the name of the sequel to the novel "The Alienist" by author Caleb Carr?
Answer: The Angel of Darkness
Query: sequel to The Alienist Caleb Carr
@@
Doc 1: Caleb Carr. THE ALIENIST. A Novel. Vintage Books Random House New York. Contents. Acknowledgments. Warning 7. Epigraph. 13. Prologue. 15. 1896. 18. 1. 21.
Doc 2: Book review: The Alienist by Caleb Carr · Comments5. The Alienist by Caleb Carr. Rating: 5 out of 5. 1896 – 1897. New York City, 1896 is a time
Doc 3: The Alienist: A Novel (2017) · The Angel of Darkness (2018) · The Lost City of Z (2019) · The Devil in the White City (2019) · A Gentleman in Moscow (2019)
Doc 4: The sequel to the acclaimed national bestseller The Alienist, Caleb Carr’s The Angel of Darkness is a breathtaking thriller set in 1897 New York.
Doc 5: The Angel of Darkness is a 1994 horror detective novel by American author Caleb Carr. It is the sequel to Carr’s best-selling debut novel
@@
Doc 1: The Alienist has been adapted for the screen, the eight-part miniseries coming to production this summer and airing in January 2018 on TNT.
Doc 2: The Alienist is a 1994 American suspense novel by Caleb Carr. Set in New York City during the winter of 1896, it follows the investigation of a
Doc 3: Imagine an alienist, an alienist is a psychiatrist in the 19th century. Okay, so who’s the alienist? John Corrigan. John Corrigan played by Daniel
Doc 4: Revisit Caleb Carr’s The Alienist by watching the classic crime drama, now streaming on HBO Max.
Doc 5: Join us this week as we explore Caleb Carr’s fictional world of monsters and monsters hunters in our conversation with Caleb about his most
@@
Appendix F Implementation Details
During experimentation, we deploy the simulation server on 4 H20 GPUs and conduct RL training on another 4 H20 GPUs. To train the simulation LLM, we conduct a lightweight SFT using Qwen-2.5-3B-Instruct, Qwen-2.5-7B-Instruct, and Qwen-2.5-14B-Instruct as the backbones. The learning rate is set to be 1e-6. To train ZeroSearch, we adopt three reinforcement learning algorithms: REINFORCE, GRPO, and PPO. For both REINFORCE and GRPO, the policy LLM is trained with a learning rate of 1e-6, and five responses are sampled per prompt. In the PPO setting, the policy LLM is trained with a learning rate of 1e-6, while the value model is trained with a separate learning rate of 1e-5. We apply Generalized Advantage Estimation (GAE) with hyperparameters $\lambda=1$ and $\gamma=1$. For Qwen-2.5-7B (Base/Instruct), the initial noise probability $p_{s}$ and final noise probability $p_{e}$ are set to 0 and 0.75. For Qwen-2.5-3B (Base/Instruct), they are set to 0 and 0.25. For LLaMA-3.2-3B (Base/Instruct), they are set to 0.25 and 0.5. Unless otherwise specified, all RL-based methods use REINFORCE as the default training algorithm. The fine-tuned Qwen-2.5-14B-Instruct model serves as the default simulated search engine across all experiments. During inference, we uniformly use Google Web Search as the search engine for all methods.
Appendix G Broad Impact
This work introduces a novel reinforcement learning (RL) training framework that enhances the search capabilities of large language models (LLMs) without requiring interaction with real-world search engines. We believe this approach can significantly reduce the cost of RL training and encourage future research that leverages LLMs as simulators of the real world, potentially yielding positive social impacts. We do not foresee any negative societal impacts arising from this work.